# Optimal Protocols for Continual Learning via Statistical Physics and Control Theory

### v1

@@ Abstract
Artificial neural networks often struggle with catastrophic forgetting when learning multiple tasks sequentially, as training on new tasks degrades the performance on previously learned ones. Recent theoretical work has addressed this issue by analysing learning curves in synthetic frameworks under predefined training protocols. However, these protocols relied on heuristics and lacked a solid theoretical foundation assessing their optimality. In this paper, we fill this gap combining exact equations for training dynamics, derived using statistical physics techniques, with optimal control methods. We apply this approach to teacher-student models for continual learning and multi-task problems, obtaining a theory for task-selection protocols maximising performance while minimising forgetting. Our theoretical analysis offers non-trivial yet interpretable strategies for mitigating catastrophic forgetting, shedding light on how optimal learning protocols can modulate established effects, such as the influence of task similarity on forgetting. Finally, we validate our theoretical findings on real-world data.
@@ 1 Introduction
Despite the significant interest in transfer learning and catastrophic forgetting, mitigation strategies considered thus far were pre-defined heuristics, offering no guarantees of optimality. In contrast, here we address the problem of identifying the optimal protocol to minimise forgetting. Specifically, we focus on replay as a prototypical mitigation strategy and employ optimal control theory to determine the optimal training protocol that maximises performance across different tasks.
@@
In this work, we combine dimensionality-reduction techniques from statistical physics (Saad & Solla, 1995a; b; Biehl & Schwarze, 1995) and Pontryagin’s maximum principle from control theory (Feldbaum, 1955; Pontryagin, 1957; Kopp, 1962) to derive optimal task-selection protocols for the training dynamics of a neural network in a continual learning setting. Pontryagin’s principle works efficiently in low-dimensional deterministic systems, hence requiring the statistical physics approach to neural networks (Engel, 2001), where the evolution of high-dimensional stochastic systems are condensed to a few key order parameters governed by ordinary differential equations (ODEs) (Saad & Solla, 1995a; b; Biehl & Schwarze, 1995). Specifically, we consider the teacher-student framework of Lee et al. (2021)—a prototype continual learning setting amenable to analytic characterisation. Our main contributions are:
@@
  • We demonstrate that the insights from our optimal strategies in synthetic settings transfer to real datasets. Specifically, we show the efficacy of our pseudo-optimal strategy on a continual learning task using the Fashion-MNIST dataset, where it optimally interpolates between simple heuristic strategies based on the problem’s parameters.
@@
In recent years, several theoretical works on online learning dynamics in one-hidden-layer neural networks have addressed a range of machine learning problems, including over-parameterisation (Goldt et al., 2019), algorithmic analysis (Refinetti et al., 2021; Srinivasan et al., 2024), and learning strategies (Lee et al., 2021; Saglietti et al., 2022; Sarao Mannelli et al., 2024). However, these studies have not explored the problem from an optimal control perspective.
Early works addressed the optimality of hyperparameters in high-dimensional online learning for committee machines via control theory. These studies focused on optimising the learning rate (Saad & Rattray, 1997; Rattray & Saad, 1998; Schlösser et al., 1999), the regularisation (Saad & Rattray, 1998), and the learning rule (Rattray & Saad, 1997). However, to the best of our knowledge, the problem of optimal task selection has not been explored yet. Carrasco-Davis et al. (2023) and Li et al. (2024) applied optimal control to the dynamics of connectivist models of behaviour, but their analysis was limited to low-dimensional settings. Urbani (2021) extended the Bellman equation to high-dimensional mean-field dynamical systems, though without considering learning processes.
@@ 2 Model-based theoretical framework
We adopt a model-based approach to investigate the supervised learning of multiple tasks. Following Lee et al. (2021; 2022), we consider a teacher-student framework (Gardner & Derrida, 1989) where a “student" neural network is trained on synthetic inputs $\bm{x}\in\mathbb{R}^{N}$, drawn i.i.d. from a standard Gaussian distribution, $x_{i}\sim\mathcal{N}(0,1)$. The labels for each task $t=1,\ldots,T$ are generated by single-layer “teacher" networks: $y^{(t)}=g_{*}(\bm{x}\cdot\bm{w}_{*}^{(t)}/\sqrt{N})$, where $\bm{W}_{*}=(\bm{w}_{*}^{(1)},\ldots,\bm{w}_{*}^{(T)})^{\top}\in\mathbb{R}^{T\times N}$ denote the corresponding teacher vectors, and $g_{*}$ the activation function. The student is a two-layer neural network with $K$ hidden units, first-layer weights $\bm{W}=\left(\bm{w}_{1},\ldots,\bm{w}_{K}\right)^{\top}\in\mathbb{R}^{K\times N}$, activation function $g$, and second-layer weights $\bm{v}\in\mathbb{R}^{K}$, that outputs the prediction:
@@
Following a standard multi-headed approach to continual learning (Zenke et al., 2017; Farquhar & Gal, 2018), we allow for task-dependent readout weights: $\bm{V}=(\bm{v}^{(1)},\ldots,\bm{v}^{(T)})^{\top}\in\mathbb{R}^{T\times K}$. While the readout is switched during training according to the task under consideration, the first-layer weights are shared across tasks. A pictorial representation of this model is displayed in Fig. 1. Training is performed via Stochastic Gradient Descent (SGD) on the squared loss of $y^{(t)}$ and $\hat{y}^{(t)}=\hat{y}(\bm{x};\bm{W},\bm{v}^{(t)})$, in the online regime, where at each training step the algorithmic update is computed using a new sample $(\bm{x},y^{(t)})$.
@@
$\varepsilon_{t}\left(\bm{W},\bm{V},\bm{W}_{*}\right)\coloneqq\frac{1}{2}\Big{<}\Big{(}y^{(t)}-\hat{y}^{(t)}\Big{)}^{2}\Big{>}=\frac{1}{2}\mathbb{E}_{\bm{x}}\left[\left(g^{*}\left(\frac{\bm{w}_{*}^{(t)}\cdot\bm{x}}{\sqrt{N}}\right)-\hat{y}(\bm{x};\bm{W},\bm{v}^{(t)})\right)^{2}\right]\,,$ (2)
@@
First, we consider the set of equations describing the training dynamics for a given task-selection protocol. As further discussed in Appendix A, in the limit of large input dimension $N\to\infty$ with $K,T\sim\mathcal{O}_{N}(1)$, the dynamics of the generalisation error is entirely captured by the evolution of the student readouts $\bm{V}$ and a set of low dimensional matrices—called overlaps in the statistical physics literature:
$M_{kt}\coloneqq\frac{\bm{w}_{k}\cdot\bm{w}_{*}^{(t)}}{N}\,,$ $Q_{kh}\coloneqq\frac{\bm{w}_{k}\cdot\bm{w}_{h}}{N}\,,$ $S_{tt^{\prime}}\coloneqq\frac{\bm{w}_{*}^{(t)}\cdot\bm{w}_{*}^{(t^{\prime})}}{N}\,,$ (3)
for all $k,h=1,\ldots,K$ and $t=1,\ldots,T$. For the remainder of the paper we consider $K=T$, to guarantee that the student network has enough capacity to learn all tasks perfectly. Teacher vectors are normalised, while the task similarity is tuned by a parameter $\gamma$, so that $S_{tt^{\prime}}=\delta_{t,t^{\prime}}+\gamma(1-\delta_{t,t^{\prime}})$. For simplicity, it is useful to encode all the dynamical degrees of freedom of interest—the overlaps and the readout weights—in the same vector. We use the shorthand notation $\mathbb{Q}=({\rm vec}(\bm{Q}),{\rm vec}(\bm{M}),{\rm vec}(\bm{V}))^{\top}\in\mathbb{R}^{K^{2}+2KT}$. Following Lee et al. (2021), we write a set of ODEs
$\frac{{\rm d}{\mathbb{Q}}(\alpha)}{{\rm d}\alpha}=f_{\mathbb{Q}}\left(\mathbb{Q}(\alpha),\bm{u}(\alpha)\right)\quad\quad\text{with }\alpha\in(0,\alpha_{F}]\;,$ (4)
$\alpha$ denoting the effective training time—i.e., the ratio between training epochs and input dimension $N$, as detailed in Appendix A—and $\bm{u}$ the dynamical variables that we want to control optimally. In particular, we study the optimal schedules for task-selection $t_{c}(\alpha)$ and learning rate $\eta(\alpha)$. Here, $t_{c}(\alpha)\in\{1\,,\ldots\,,T\}$ indicates on which task the student is trained at time $\alpha$. The specific form of the functions $f_{\mathbb{Q}}$ is derived in Appendix A. The initial condition $\mathbb{Q}(0)$ matches the initialisation of the SGD algorithm. In particular, the initial first-layer weights and readout weights are drawn i.i.d. from a normal distribution with variances of $10^{-3}$ and $10^{-2}$, respectively. Notably, the trajectory appears to be largely independent of the specific initialisation of the first-layer weights. For instance, in Fig. 2, simulations and theory correspond to different initialisations, yet the curves show excellent agreement. Let us stress that Eq. 4 is a low-dimensional deterministic equation that fully captures the high-dimensional stochastic dynamics of SGD as $N\to\infty$. This dimensionality reduction is crucial to apply the optimal control techniques presented in the next section.
@@
Our first main contribution is to derive training strategies that are “optimal” with respect to the generalisation performance at the end of training and on all tasks. In practice, the goal of the optimisation process is to minimise a linear combination of the generalisation errors on the different tasks at the final training time $\alpha_{F}$:
@@
where the coefficients $c_{t}$ identify the relative importance of different tasks and $\varepsilon_{t}$ denotes the infinite-dimensional limit of the average generalisation error on task $t$, as defined in Eq. 2. Crucially, we have an analytic expression for $\varepsilon_{t}$, derived in Appendix A. In the remainder of the paper, we assume equally important tasks $c_{t}=1/T$. As customary in optimal control theory (Pontryagin, 1957), we adopt a variational approach to solve the problem. We define the cost function
@@
where the conjugate order parameters ^$\hat{\mathbb{Q}}=({\rm vec}(\hat{\bm{Q}}),{\rm vec}(\hat{\bm{M}}),{\rm vec}(\hat{\bm{V}}))^{\top}$ enforce the training dynamics in the training interval $\alpha\in[0,\alpha_{F}]$. Finding the optimal protocol amounts to minimising the cost function $\mathcal{F}$ with respect to $\mathbb{Q}$, $\hat{\mathbb{Q}}$, and $\bm{u}$. We defer the details of this variational procedure to Appendix A, and present only the main steps here. The minimisation with respect to $\mathbb{Q}$ provides a set of equations for the backward dynamics of the conjugate parameters
@@
where $\mathcal{U}$ is the set of allowed controls, that can be either continuous or discrete. For instance, for task selection we take ${u}(\alpha)=t_{c}(\alpha)$ and $\mathcal{U}=\{1\,,2\,\ldots\,,T\}$, where we use the notation $t_{c}(\alpha)$ to indicate the current task, i.e., the task on which the student is trained at time $\alpha$. When optimising over both task selection and learning rate schedule we take $\bm{u}=(t_{c},\eta)$ and $\mathcal{U}=\{1\,,2\,\ldots\,,T\}\times\mathbb{R}^{+}$. Crucially, the optimal control equations 4, 7, and 9 must be iterated until convergence, starting from an initial guess on $\bm{u}$. Let us stress that the space $\mathcal{U}$ of possible controls is high-dimensional and hence it is not feasible to explore it via greedy search strategies.
@@ 3 Results and applications
The theoretical framework of Sec. 2 is extremely flexible and can be applied to a variety of settings. We will present a technical paper focused on the method in a specialised venue, while in this section we focus on a particular setting and investigate in detail the impact of optimal training protocols on both synthetic and real tasks. First, we consider the synthetic teacher-student framework. We compute optimal task-selection protocols, investigating their structure and interplay with task similarity and learning rate schedule. Then, we transfer the insights gained from the interpretation of the optimal strategies in synthetic settings to applications on real datasets.

### v2

@@ Abstract
Artificial neural networks often struggle with catastrophic forgetting when learning multiple tasks sequentially, as training on new tasks degrades the performance on previously learned tasks. Recent theoretical work has addressed this issue by analysing learning curves in synthetic frameworks under predefined training protocols. However, these protocols relied on heuristics and lacked a solid theoretical foundation assessing their optimality. In this paper, we fill this gap by combining exact equations for training dynamics, derived using statistical physics techniques, with optimal control methods. We apply this approach to teacher-student models for continual learning and multi-task problems, obtaining a theory for task-selection protocols maximising performance while minimising forgetting. Our theoretical analysis offers non-trivial yet interpretable strategies for mitigating catastrophic forgetting, shedding light on how optimal learning protocols modulate established effects, such as the influence of task similarity on forgetting. Finally, we validate our theoretical findings with experiments on real-world data.
@@ 1 Introduction
Despite the significant interest in transfer learning and catastrophic forgetting, mitigation strategies considered thus far were pre-defined heuristics, offering no guarantees of optimality. In contrast, here we address the problem of identifying the optimal protocol to minimise forgetting. Specifically, we focus on replay as a prototypical mitigation strategy. We use optimal control theory to determine the optimal training protocol that maximises performance across different tasks.
@@
In this work, we combine dimensionality-reduction techniques from statistical physics (Saad & Solla, 1995a; b; Biehl & Schwarze, 1995) and Pontryagin’s maximum principle from control theory (Feldbaum, 1955; Pontryagin, 1957; Kopp, 1962). This approach allows us to derive optimal task-selection protocols for the training dynamics of a neural network in a continual learning setting. Pontryagin’s principle works efficiently in low-dimensional deterministic systems. Hence, applying it to neural networks requires the statistical physics approach (Engel, 2001), which reduces the evolution of high-dimensional stochastic systems to a few key order parameters governed by ordinary differential equations (ODEs) (Saad & Solla, 1995a; b; Biehl & Schwarze, 1995). Specifically, we consider the teacher-student framework of Lee et al. (2021)—a prototype continual learning setting amenable to analytic characterisation. Our main contributions are:
@@
  • We demonstrate that the insights from our optimal strategies in synthetic settings transfer to real datasets. Specifically, we show the efficacy of our pseudo-optimal strategy on a continual learning task using the Fashion-MNIST dataset. Here, the pseudo-optimal strategy effectively interpolates between simple heuristics depending on the problem’s parameters.
@@
Recent theoretical works on online dynamics in one-hidden-layer neural networks have addressed various learning problems, including over-parameterisation (Goldt et al., 2019), algorithmic analysis (Refinetti et al., 2021; Srinivasan et al., 2024), and learning strategies (Lee et al., 2021; Saglietti et al., 2022; Sarao Mannelli et al., 2024). However, these studies have not explored the problem from an optimal control perspective.
Early works addressed the optimality of hyperparameters in high-dimensional online learning for committee machines via control theory. These studies focused on optimising the learning rate (Saad & Rattray, 1997; Rattray & Saad, 1998; Schlösser et al., 1999), the regularisation (Saad & Rattray, 1998), and the learning rule (Rattray & Saad, 1997). However, to the best of our knowledge, the problem of optimal task selection has not been explored yet. Carrasco-Davis et al. (2023) and Li et al. (2024) applied optimal control to the dynamics of connectionist models of behaviour, but their analysis was limited to low-dimensional and finite-dimensional settings. Urbani (2021) extended the Bellman equation to high-dimensional mean-field dynamical systems, though without considering learning processes.
@@ 2 Model-based theoretical framework
We adopt a model-based approach to investigate the supervised learning of multiple tasks. Following Lee et al. (2021; 2022), we consider a teacher-student framework (Gardner & Derrida, 1989). A “student" neural network is trained on synthetic inputs $\bm{x}\in\mathbb{R}^{N}$, drawn i.i.d. from a standard Gaussian distribution, $x_{i}\sim\mathcal{N}(0,1)$. The labels for each task $t=1,\ldots,T$ are generated by single-layer “teacher" networks: $y^{(t)}=g_{*}(\bm{x}\cdot\bm{w}_{*}^{(t)}/\sqrt{N})$, where $\bm{W}_{*}=(\bm{w}_{*}^{(1)},\ldots,\bm{w}_{*}^{(T)})^{\top}\in\mathbb{R}^{T\times N}$ denote the corresponding teacher vectors, and $g_{*}$ the activation function. The student is a two-layer neural network with $K$ hidden units, first-layer weights $\bm{W}=\left(\bm{w}_{1},\ldots,\bm{w}_{K}\right)^{\top}\in\mathbb{R}^{K\times N}$, activation function $g$, and second-layer weights $\bm{v}\in\mathbb{R}^{K}$. It outputs the prediction:
@@
Following a standard multi-headed approach to continual learning (Zenke et al., 2017; Farquhar & Gal, 2018), we allow for task-dependent readout weights: $\bm{V}=(\bm{v}^{(1)},\ldots,\bm{v}^{(T)})^{\top}\in\mathbb{R}^{T\times K}$. Specifically, the readout for task $t$ is updated only when that task is presented. While the readout is switched during training according to the task under consideration, the first-layer weights are shared across tasks. A pictorial representation of this model is displayed in Fig. 1. Training is performed via Stochastic Gradient Descent (SGD) on the squared loss of $y^{(t)}$ and $\hat{y}^{(t)}=\hat{y}(\bm{x};\bm{W},\bm{v}^{(t)})$. We consider the online regime, where at each training step the algorithmic update is computed using a new sample $(\bm{x},y^{(t)})$.
@@
$\varepsilon_{t}\left(\bm{W},\bm{V},\bm{W}_{*}\right)\coloneqq\frac{1}{2}\Big{<}\Big{(}y^{(t)}-\hat{y}^{(t)}\Big{)}^{2}\Big{>}=\frac{1}{2}\mathbb{E}_{\bm{x}}\left[\left(g^{*}\left(\frac{\bm{w}_{*}^{(t)}\cdot\bm{x}}{\sqrt{N}}\right)-\hat{y}(\bm{x};\bm{W},\bm{v}^{(t)})\right)^{2}\right]\,.$ (2)
@@
$\lambda_{k}\coloneqq\frac{\bm{x}\cdot\bm{w}_{k}}{\sqrt{N}}\;,\;\;k=1,\ldots,K\,,$ and $\lambda_{*}^{(t)}\coloneqq\frac{\bm{x}\cdot\bm{w}^{(t)}_{*}}{\sqrt{N}}\;,\;\;t=1,\ldots,T.$ (3)
Eq. 3 defines jointly Gaussian variables with zero mean and second moments given by
$\begin{split}M_{kt}&\coloneqq\mathbb{E}_{\bm{x}}\left[\lambda_{k}\lambda_{*}^{(t)}\right]=\frac{\bm{w}_{k}\cdot\bm{w}_{*}^{(t)}}{N}\;,\\ Q_{kh}&\coloneqq\mathbb{E}_{\bm{x}}\left[\lambda_{k}\lambda_{h}\right]=\frac{\bm{w}_{k}\cdot\bm{w}_{h}}{N}\;,\\ S_{tt^{\prime}}&\coloneqq\mathbb{E}_{\bm{x}}\left[\lambda_{*}^{(t)}\lambda_{*}^{(t^{\prime})}\right]=\frac{\bm{w}_{*}^{(t^{\prime})}\cdot\bm{w}_{*}^{(t)}}{N}\;,\end{split}$ (4)
called overlaps in the statistical physics literature. Therefore, the dynamics of the generalisation error is entirely captured by the evolution of the student readouts $\bm{V}$ and the overlaps. As shown in Lee et al. (2021; 2022), we can track the evolution of the generalisation error in the high-dimensional. We leverage this description and optimal control theory to derive optimal training protocols for multi-task learning. In particular, we optimise over task selection and learning rate.
@@
First, we derive equations governing the dynamics of the overlaps and readouts under a given task-selection protocol. These equations fully determine the evolution of the generalization error. For the remainder of the paper, we consider $K=T$ to guarantee that the student network has enough capacity to learn all tasks perfectly. Teacher vectors are normalised, and the task similarity is tuned by a parameter $\gamma$, so that $S_{tt^{\prime}}=\delta_{t,t^{\prime}}+\gamma(1-\delta_{t,t^{\prime}})$. For simplicity, it is useful to encode all the relevant degrees of freedom—namely, the overlaps and the readout weights—in the same vector. We use the shorthand notation $\mathbb{Q}=({\rm vec}(\bm{Q}),{\rm vec}(\bm{M}),{\rm vec}(\bm{V}))^{\top}\in\mathbb{R}^{K^{2}+2KT}$. As further discussed in Appendix A, in the limit of large input dimension $N\to\infty$ with $K,T\sim\mathcal{O}_{N}(1)$, the training dynamics is described by a set of ODEs
$\frac{{\rm d}{\mathbb{Q}}(\alpha)}{{\rm d}\alpha}=f_{\mathbb{Q}}\left(\mathbb{Q}(\alpha),\bm{u}(\alpha)\right)\quad\quad\text{with }\alpha\in(0,\alpha_{F}]\;.$ (5)
The parameter $\alpha$ denotes the effective training time—the ratio between training steps and input dimension $N$. The vector $\bm{u}$ encodes the dynamical variables that we want to control optimally. In particular, we study the optimal schedules for task-selection $t_{c}(\alpha)$ and learning rate $\eta(\alpha)$. Here, $t_{c}(\alpha)\in\{1\,,\ldots\,,T\}$ indicates on which task the student is trained at time $\alpha$. The specific form of the functions $f_{\mathbb{Q}}$ is derived in Appendix A. The initial condition $\mathbb{Q}(0)$ matches the initialisation of the SGD algorithm. In particular, the initial first-layer weights and readout weights are drawn i.i.d. from a normal distribution with variances of $10^{-3}$ and $10^{-2}$, respectively. Notably, the trajectory appears to be largely independent of the specific initialisation of the first-layer weights. For instance, in Fig. 2, simulations and theory correspond to different initialisations, yet the curves show excellent agreement. Let us stress that Eq. 5 is a low-dimensional deterministic equation that fully captures the high-dimensional stochastic dynamics of SGD as $N\to\infty$. This dimensionality reduction is crucial to apply the optimal control techniques presented in the next section.
@@
Our first main contribution is to derive training strategies that are optimal with respect to the generalisation performance at the end of training and on all tasks. In practice, the goal of the optimisation process is to minimise a linear combination of the generalisation errors on the different tasks at the final training time $\alpha_{F}$:
@@
where the coefficients $c_{t}$ identify the relative importance of different tasks and $\varepsilon_{t}$ denotes the infinite-dimensional limit of the average generalisation error on task $t$, as defined in Eq. 2. Crucially, we have an analytic expression for $\varepsilon_{t}$, derived in Appendix A. In the remainder of the paper, we assume equally important tasks $c_{t}=1/T$. As customary in optimal control theory (Pontryagin, 1957), we adopt a variational approach to solve the problem. We define the cost functional
@@
where the conjugate order parameters ^$\hat{\mathbb{Q}}=({\rm vec}(\hat{\bm{Q}}),{\rm vec}(\hat{\bm{M}}),{\rm vec}(\hat{\bm{V}}))^{\top}$ enforce the training dynamics in the training interval $\alpha\in[0,\alpha_{F}]$. Finding the optimal protocol amounts to minimising the cost functional $\mathcal{F}$ with respect to $\mathbb{Q}$, $\hat{\mathbb{Q}}$, and $\bm{u}$. We defer the details of this variational procedure to Appendix A, and present only the main steps here. For a general introduction to the control methods adopted here, see, e.g., Kirk (2004). The minimisation with respect to $\mathbb{Q}$ provides a set of equations for the backward dynamics of the conjugate parameters
@@
where $\mathcal{U}$ is the set of allowed controls, that can be either continuous or discrete. For instance, for task selection we take ${u}(\alpha)=t_{c}(\alpha)$ and $\mathcal{U}=\{1\,,2\,\ldots\,,T\}$, where we use the notation $t_{c}(\alpha)$ to indicate the current task, i.e., the task on which the student is trained at time $\alpha$. When optimising over both task selection and learning rate schedule we take $\bm{u}=(t_{c},\eta)$ and $\mathcal{U}=\{1\,,2\,\ldots\,,T\}\times\mathbb{R}^{+}$. Crucially, the optimal control equations 5, 8, and 10 must be iterated until convergence, starting from an initial guess on $\bm{u}$, which can be for instance taken at random. Let us stress that the space $\mathcal{U}$ of possible controls is high-dimensional and hence it is not feasible to explore it via greedy search strategies.
@@ 3 Results and applications
The theoretical framework of Sec. 2 is extremely flexible and can be applied to a variety of settings. We will present a technical paper focused on the method in a specialised venue. In this section, we focus on specific settings and investigate in detail the impact of optimal training protocols on both synthetic and real tasks. First, we consider the synthetic teacher-student framework. We compute optimal task-selection protocols, investigating their structure and interplay with task similarity and learning rate schedule. Then, we transfer the insights gained from the interpretation of the optimal strategies in synthetic settings to applications on real datasets.